trying to help my son on this and struggling. Solve for X
5(x-3) = 3 (x-5)
I am not sure how to isolate one X on one side of the equation?

Answer by Alan3354(69443)

5(x-3) = 3 (x-5)
5x - 15 = 3x - 15
5x = 3x
2x = 0
x = 0